Please help me solve: Susan Marciano invested part of her $18,000 bonus in a fund that paid an 8% profit and invested the rest in stock that suffered a 3% loss. Find the amount of each investment if her overall net profit was $780.
They want to know: The amount invested at 8% is:
and the amount invested in stock is:
: Please help me solve: Susan Marciano invested part of her $18,000 bonus in a fund that paid an 8% profit and invested the rest in stock that suffered a 3% loss. Find the amount of each investment if her overall net profit was $780.
They want to know: The amount invested at 8% is:
and the amount invested in stock is:

Answer by orca(336)

Let x be the amount invested in funds that earned 8%.
Then the amount invested in stock that suffered 3% loss is 18,000 - x.
The profit from investment in funds is 8%x
The loss from investment in stock is 3%(18,000 - x).
The net profit is 8%x - 3%(18,000 - x).
As the net profit given in the problem is $780, we have:
8%x - 3%(18,000 - x) = 780
Solving for x, we have:
8x - 3(18,000 - x) = 78000 (multiply both sides by 100)
8x - 54000 + 3x = 78000
11x = 132000
x = 132000/11
x = 12,000
So
the amount invested in funds that earned 8% is 12,000.
the amount invested in stock that suffered 3% loss is 18,000 - x = 18000 - 12000 = 6,000.